# Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition

### v1

@@ Abstract
Layer removal has emerged as a promising approach for compressing large language models (LLMs) by leveraging redundancy within layers to reduce model size and accelerate inference. However, this technique often compromises internal consistency, leading to performance degradation and instability, with varying impacts across different model architectures. In this work, we propose Taco-SVD, a task-aware framework that retains task-critical singular value directions, preserving internal consistency while enabling efficient compression. Unlike direct layer removal, Taco-SVD preserves task-critical transformations to mitigate performance degradation. By leveraging gradient-based attribution methods, Taco-SVD aligns singular values with downstream task objectives. Extensive evaluations demonstrate that Taco-SVD outperforms existing methods in perplexity and task performance across different architectures while ensuring minimal computational overhead.
@@
Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023). However, their massive parameter sizes create significant computational and memory challenges, making them difficult to deploy on resource-constrained devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
@@
Figure 1: Zero-shot performance of Taco-SVD and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
@@
Despite its simplicity and computational efficiency, pure layer removal introduces two notable challenges, as shown in Figure 1. First, while task accuracy may remain stable in certain cases, perplexity (PPL) often degrades significantly, reflecting disruptions to the model’s internal consistency, which stems from impaired information flow and misaligned intermediate representations (Liang et al., 2024). This increase in PPL suggests that layer removal may affect critical transformations responsible for maintaining coherent representations, potentially reducing the model’s generalization ability on complex or unseen tasks.
Second, the impact of layer removal varies significantly across model architectures. For instance, under the same pruning ratio, LLaMA2 (Touvron et al., 2023) maintains reasonable performance, whereas LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This observation indicates that the tolerance to layer pruning is architecture-dependent. Such variability underscores a lack of robustness in pure layer removal, where certain task-critical transformations may be unevenly affected, resulting in performance inconsistencies across different architectures.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in maintaining both internal consistency and task-specific performance. While existing methods, such as LoRA-based fine-tuning (Hu et al., 2022) or lightweight layer replacements (Chen et al., 2024), aim to recover or approximate the lost functionality, they often incur substantial computational costs and do not fully address the challenge of preserving task-critical transformations in a resource-efficient manner.
To address these challenges, we propose Taco-SVD (Task-Aware Singular Value Decomposition for LLM Compression), a novel framework designed to address the limitations of pure layer removal by preserving task-critical transformations. Taco-SVD fundamentally rethinks layer removal by integrating task-aware optimization into singular value decomposition, ensuring that compression decisions align with task-specific objectives. Unlike conventional SVD, which focuses on global matrix reconstruction, Taco-SVD leverages gradient attribution to directly quantify the task-specific importance of singular values. Using a small calibration dataset from a general-purpose, unlabeled source, our method identifies and preserves the most impactful transformations within redundant layers. This approach minimizes the indiscriminate loss of critical task-specific transformations, maintaining internal consistency and representation flow while achieving efficient compression.
@@
To validate our approach, we conducted experiments on tasks such as language generation, commonsense reasoning, and mathematical reasoning using LLaMA and Mistral models. Taco-SVD achieves parameter reduction comparable to pure layer removal while improving perplexity and task performance. Notably, Taco-SVD demonstrates consistent robustness across diverse model architectures, maintaining strong performance without additional training. It also achieves strong performance with minimal calibration data—only 64 samples from a general-purpose, unlabeled dataset, demonstrating its effectiveness. Moreover, fine-tuning further enhances its performance, highlighting Taco-SVD’s robustness and efficiency in compressing large language models.
@@
Figure 2: Overview of Taco-SVD. The figure illustrates how Taco-SVD uses gradient attribution to identify task-critical singular values and applies SVD to compress redundant layers while preserving task-relevant transformations.
@@ 2 Related Work
2.1 Layer Removal Techniques and Their Impact on Model Consistency
@@
Although effective in reducing model size, layer removal often disrupts internal consistency, leading to degraded task performance and increased perplexity, as shown in our analysis building on the concept introduced by Liang et al. (2024). Additionally, its sensitivity to different architectures limits generalizability, with some models tolerating pruning better than others. These challenges highlight the need for approaches that preserve task-critical information while ensuring robust compression across diverse scenarios.
2.2 Post-Pruning Compensation and Training-Based Recovery Approaches
To mitigate the degradation caused by layer removal, post-pruning compensation methods have been developed. (Kim et al., 2024) introduced Shortened LLaMA, which incorporates LoRA (Hu et al., 2022) to recover the ability of layer removal models. However, LoRA alters the original singular value spectrum, which may diminish the contribution of pre-trained features (Shuttleworth et al., 2024). To address this issue, LLM-Streamline (Chen et al., 2024) trains a lightweight module such as feed-forward neural network (FFN) or a transformer layer to approximate the functionality of the pruned redundant layers. Although effective to some extent, these methods incur significant computational and data costs, limiting their practicality in resource-constrained scenarios.
2.3 Singular Value Decomposition for LLM Compression
Singular Value Decomposition (SVD) is a foundational technique for compressing dense models by decomposing weight matrices into low-rank structures. However, vanilla SVD focuses solely on global matrix reconstruction, providing general approximations without considering task-specific characteristics. Recent methods have enhanced SVD to better align with model and task requirements. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a calibration dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
Despite these advancements, existing methods mainly focus on optimizing compression loss without directly addressing the relationship between singular values and task-specific losses. Taco-SVD bridges this gap by integrating task-aware optimization into SVD, leveraging gradient attribution to retain task-critical singular values and align compressed models with task objectives.
@@ 3.1 Overview and Motivation
In this paper, we propose a novel task-aware compression framework named Taco-SVD, designed to compress redundant layers in large language models (LLMs) while preserving task-critical transformations. Unlike traditional layer removal methods that directly eliminate redundant components, Taco-SVD integrates task-aware singular value decomposition into the layer removal paradigm. By introducing gradient-based attributions, it systematically aligns the importance of singular values with task-specific objective, ensuring that the most essential components within redundant layers are retained during truncation.
Figure 2 provides an overview of Taco-SVD. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) applying task-aware singular value decomposition (SVD) to retain the most critical singular values and their corresponding singular vectors. Below, we describe each step in detail.
@@ 3.2 Redundant Layer Selection
The first step in Taco-SVD is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
@@
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, Taco-SVD compresses their weight matrices using a task-aware approach to preserve task-relevant transformations.
3.3 Task-Aware Singular Value Decomposition
3.3.1 Motivation for Task-Aware SVD
Singular value decomposition (SVD) is a standard technique for compressing weight matrices by providing a global low-rank approximation. Traditional SVD methods retain the top-$k$ singular values based solely on their magnitudes, which are task-agnostic and often fail to capture task-relevant components. This can lead to the removal of critical transformations necessary for downstream tasks (Hsu et al., 2022; Wang et al., 2024).
To address this limitation, Taco-SVD introduces task awareness into SVD by using gradient-based attribution to evaluate the task-specific importance of each singular value and its corresponding singular vectors. This ensures that only the most task-relevant transformations are preserved.
@@ 3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W\in\mathbb{R}^{m\times n}$, SVD decomposes $W$ as:
$W=U\Sigma V^{T}$ (2)
@@
3.3.3 Task-Aware Singular Value Importance
To assess the task-specific importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Task-Aware Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the task loss $\mathcal{L}_{\text{task}}$. Using its first-order Taylor expansion, the alteration in the task loss $L$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (3)
where $\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the task loss, and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{task}}$ with respect to $W$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{task}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (4)
where $G=\frac{\partial\mathcal{L}_{\text{task}}}{\partial W}$ is the gradient of the task loss with respect to $W$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
@@ 3.3.3 Task-Aware Singular Value Importance
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to task performance, considering both its magnitude and alignment with the task gradient. Larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of task relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{task}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (6)
This ensures that $W_{\text{task}}$ preserves the transformations most critical to task performance, achieving efficient compression while maintaining task consistency.
@@ 3.4 Layer-wise Low-Rank Decomposition
Taco-SVD applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Task-Aware Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves task-relevant transformations.
@@ 4.1 Experimental Setup
To validate the effectiveness of Taco-SVD, we conducted experiments across multiple large language models (LLMs) and datasets, focusing on evaluating task performance, cross-model robustness, and computational efficiency. Below, we detail the models, datasets, baselines, and experimental configurations.
@@
We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral-7b (Jiang et al., 2023). These models enable a comprehensive assessment of our method across varying sizes and architectures, highlighting its generability and robustness.
Datasets
Ten datasets are used as evaluation benchmarks including three language generation datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), C4 (Raffel et al., 2020) and seven datasets for common sense and mathematical reasoning include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018) for common sense reasoning, and MathQA (Amini et al., 2019) for mathematical reasoning.
Metric
We use perplexity and accuracy for language generation tasks and reasoning tasks, respectively. All reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024).
Baselines
We compare Taco-SVD against training-free and training-based compression methods at the same compression ratio. For training-free methods, we select the layer removal technique LaCo (Yang et al., 2024), as well as a structured pruning method SliceGPT (Ashkboos et al., 2024) and the SVD-based LLM compression method ASVD (Yuan et al., 2023). For training-based methods, we compare Taco-SVD with LoRA (Hu et al., 2022) and the layer replacement method LLM-Streamline (Chen et al., 2024) with two variants: LLM-Streamline-FFN and LLM-Streamline-Layer, which replace redundant layers with Feed-Forward Neural Network (FFN) and Transformer Layer (Layer), respectively.

### v2

# Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression

@@ Abstract
Layer removal is an effective technique for compressing large language models (LLMs) by reducing redundancy and improving inference efficiency. However, indiscriminate pruning disrupts representation stability, leading to performance degradation. We propose GRASP (Gradient-based Retention of Adaptive Singular Parameters), which preserves representation-critical singular values to mitigate these effects. Unlike direct layer removal, GRASP leverages gradient-based attribution on a syntax- and semantics-rich dataset to guide the selection of representation-critical singular values. By selectively applying singular value decomposition (SVD) to affected layers, GRASP achieves efficient compression while maintaining representation stability with minimal overhead. Experiments across multiple LLMs show that GRASP consistently outperforms existing compression methods in perplexity and downstream task performance. The source code is available at GRASP.
@@
Rethinking Layer Removal: A Hybrid Pruning Framework Combining Layer Removal and Singular Value Selection for Efficient LLM Compression
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, including language generation, reasoning, and question answering (Brown et al., 2020; Touvron et al., 2023b). However, their massive parameter sizes pose computational and memory challenges, hindering deployment on resource-limited devices (Zhou et al., 2024). To address this, model compression techniques such as quantization (Frantar et al., 2022; Lin et al., 2024; Xiao et al., 2023), knowledge distillation (Gu et al., 2023; Xu et al., 2024), and pruning(Sun et al., 2024; Ashkboos et al., 2024) have been widely explored. Among these, structured pruning methods, especially layer removal (Men et al., 2024; Yang et al., 2024; Song et al., 2024), have shown promise by leveraging redundancy within layers of LLMs to achieve significant parameter reductions and inference speedup.
@@
Figure 1: Zero-shot performance of GRASP and pure layer removal under 20% compression ratio. Dashed lines represent perplexity on WikiText-2, while bar charts show the percentage of the original model’s average accuracy on seven reasoning datasets.
@@
Despite its simplicity and efficiency, layer removal introduces two main challenges, as shown in Figure 1. First, while task accuracy may remain stable in some cases, the perplexity (PPL) often degrades significantly, reflecting disruptions in the model’s internal representation consistency. This degradation arises from impaired information flow and misaligned intermediate representations (Liang et al., 2024). The sharp increase in PPL suggests that layer removal eliminates critical transformations necessary for maintaining coherent representation structures, potentially limiting the model’s generalization on complex or unseen tasks. Second, the impact of layer removal varies significantly across models. For example, under the same pruning ratio, LLaMA2 (Touvron et al., 2023b) maintains reasonable performance, while LLaMA3 (Dubey et al., 2024) experiences a steep performance drop. This discrepancy suggests that layer removal affects models unevenly, as certain structural transformations are more essential for maintaining stable representations, leading to performance inconsistencies across models.
Our analysis suggests that these issues arise from the removal of transformations that play an essential role in preserving internal consistency and maintaining structured representations. In this paper, we introduce GRASP (Gradient-based Retention of Adaptive Singular Parameters), a novel hybrid pruning framework that integrates layer removal with gradient-based singular value selection for efficient LLM compression. Unlike conventional pruning techniques that either remove full layers or apply global SVD compression, GRASP selectively applies singular value decomposition (SVD) only to layers affected by pruning. A key component of GRASP is its use of a small, syntax- and semantics-rich dataset (e.g., WikiText-2), which serves as a representation evaluation dataset to guide the selection of singular values. Instead of relying solely on magnitude-based ranking, GRASP leverages gradient attribution computed from this dataset to dynamically identify and retain singular values critical for preserving language representations. This ensures that despite layer removal, the model retains essential syntactic and semantic structures, minimizing degradation in perplexity and downstream task performance.
@@
We evaluate GRASP on 10 datasets and five models from two distinct LLM families (LLaMA, Mistral). Extensive experiments validate the effectiveness of GRASP. It achieves compression rates comparable to pure layer removal while significantly improving perplexity and task performance across a wide rage of downstream tasks. Notably, GRASP maintains strong robustness across different models without requiring additional training. Furthermore, it achieves strong performance with minimal calibration data—using only 64 samples from a general-purpose, unlabeled dataset, demonstrating the efficiency of our approach. Additionally, post-training compensation further enhances GRASP’s performance, highlighting its adaptability in compressing large language models while preserving their linguistic capabilities.
@@
Figure 2: (A): Overview of GRASP. The figure illustrates how GRASP uses gradient attribution to identify representation-critical singular values and applies SVD to compress redundant layers while preserving representation-relevant transformations. (B): SVD-based LLM compression and (C): Layer removal compression methods
@@ 2 Related Work
2.1 Layer Removal-based LLM Compression
@@
Although effective in reducing model size, layer removal disrupts representation coherence, leading to performance degradation and increased perplexity, as analyzed by Liang et al. (2024). To mitigate this, post-training compensation methods have been proposed. Kim et al., 2024 introduced Shortened LLaMA, which employs LoRA (Hu et al., 2022) to restore pruned models’ capabilities. However, LoRA modifies the singular value spectrum, potentially weakening pre-trained features (Shuttleworth et al., 2024). LLM-Streamline (Chen et al., 2024) addresses this by training a lightweight module, such as an FFN or transformer layer, to approximate the removed layers. While effective, these methods impose high computational and data costs, limiting feasibility in resource-constrained settings.
2.2 SVD-based LLM Compression
Singular Value Decomposition (SVD) is widely used in model compression to decompose weight matrices into low-rank structures, typically selecting top-k singular values based on Frobenius norm reconstruction loss. Recent methods have enhanced SVD to reduce LLM compression error. FWSVD (Hsu et al., 2022) incorporates Fisher information to reweight the importance of parameters before applying SVD. ASVD (Yuan et al., 2023) uses activation patterns from a representation evaluation dataset to scale weight matrices, reducing compression-induced activation errors. SVD-LLM (Wang et al., 2024) applies truncation-aware data whitening and layer-wise updates to ensure a direct relationship between singular values and compression loss. Additionally, (Yu and Wu, 2023; Chavan et al., 2024; Ji et al., 2024) present another paradigm for low-rank compression of LLMs, where eigenvalue decomposition is applied to output activations, approximating the output features of each layer with low-rank matrices.
GRASP differs from these methods in three key aspects. First, rather than minimizing F2 norm reconstruction loss, GRASP optimizes a representation learning attribution loss, computed on a small representation evaluation dataset to ensure selection aligns with model representation rather than weight approximation. Second, instead of using a top-k singular value selection strategy, GRASP leverages gradient attribution to dynamically retain representation-critical singular values, which are not necessarily the largest in magnitude. Third, while SVD-based pruning is typically applied to all layers, GRASP is selectively used only in pruned layers, ensuring that layer removal does not disrupt critical transformations.
@@ 3.1 Overview and Motivation
In this paper, we propose GRASP, a novel hybrid gradient-guided compression framework designed to compress redundant layers in LLMs while preserving representation-critical transformations. Figure 2 provides an overview of GRASP. The method consists of two main steps: (1) identifying redundant layers based on cosine similarity, and (2) selecting and retaining representation-critical singular values and their corresponding singular vectors using gradient-based attribution. Below, we describe each step in detail.
@@ 3.2 Redundant Layer Selection
The first step in GRASP is identifying redundant layers. These are layers that contribute minimally to the transformation of hidden states, exhibiting high redundancy and limited impact on overall model performance. Following prior works (Song et al., 2024; Chen et al., 2024), we use cosine similarity to quantify the degree of transformation in a given layer.
@@
A high cosine similarity indicates minimal transformation, suggesting that the layer is redundant. Instead of directly removing these layers, GRASP compresses weight matrices with a gradient-based approach to retain critical internal transformations.
3.3 Gradient-Guided Singular Value Selection
3.3.1 Representation Learning Attribution Loss (RLAL)
To address this limitation, GRASP introduces gradient-based attribution to evaluate the importance of each singular value based on its contribution to model performance, rather than relying on magnitude alone. Unlike traditional SVD-based methods that select top-$k$ singular values using a reconstruction loss measured by the Frobenius norm ($F_{2}$ norm), GRASP utilizes a Representation Learning Attribution Loss (RLAL), designed specifically for general representation capability modeling.
RLAL is computed on a small, unlabeled representation evaluation dataset derived from corpora rich in syntactic and semantic structures, such as WikiText-2 and similar datasets. These datasets ensure that the retained singular values are those most crucial for maintaining the model’s ability to capture fundamental language structures, including syntax and semantics, even after compression. Unlike task-specific objectives, RLAL focuses on preserving linguistic consistency and general-purpose representation learning rather than optimizing for any particular downstream application.
Mathematically, RLAL is formulated as:
$\mathcal{L}_{\text{RLAL}}=-\sum_{t}\log P(y_{t}\mid x_{\leq t})$ (2)
where $y_{t}$ is the target token at time step $t$, $x_{\leq t}$ denotes the sequence of preceding tokens, and $P(y_{t}\mid x_{\leq t})$ represents the predicted probability of the next token given the context. This loss function aligns with the standard language modeling loss, ensuring that the model learns meaningful representations.
@@ 3.3.2 Singular Value Decomposition (SVD)
For a redundant layer’s weight matrix $W_{\text{redundant}}\in\mathbb{R}^{m\times n}$, SVD decomposes $W_{\text{redundant}}$ as:
$W_{\text{redundant}}=U\Sigma V^{T}$ (3)
@@
3.3.3 Representation-Critical Singular Value Importance
To assess the representation-critical importance of each singular value, we introduce $\Phi(\sigma_{i})$, the Representation-Critical Singular Value Importance, which quantifies the contribution of $\sigma_{i}$ to the $\mathcal{L}_{\text{RLAL}}$. Using its first-order Taylor expansion, the alteration in the $\mathcal{L}_{\text{RLAL}}$ caused by $\sigma_{i}$ is approximated as:
$\Delta\mathcal{L}_{\text{RLAL}}(\sigma_{i})\approx\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}\Delta\sigma_{i}$ (4)
where $\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}$ represents the gradient sensitivity of $\sigma_{i}$ to the representation learning attribution loss , and $\Delta\sigma_{i}$ is the perturbation applied to $\sigma_{i}$.
The gradient sensitivity is calculated by projecting the gradient of $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$ onto the singular directions:
$\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial\sigma_{i}}=\mathbf{u}_{i}^{T}G\mathbf{v}_{i}$ (5)
where $G=\frac{\partial\mathcal{L}_{\text{RLAL}}}{\partial W}$ is the gradient of the $\mathcal{L}_{\text{RLAL}}$ with respect to $W_{\text{redundant}}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ are the left and right singular vectors corresponding to $\sigma_{i}$. The detailed proof can be found in the Appendix A.1.
@@ 3.3.3 Representation-Critical Singular Value Importance
where $\Phi(\sigma_{i})$ measures the overall contribution of $\sigma_{i}$ to representation performance, considering both its magnitude and alignment with the gradient of the general-purpose model loss. A larger value of $\Phi(\sigma_{i})$ indicates greater importance.
We rank singular values in descending order of representation relevance based on $\Phi(\sigma_{i})$. Only the top-$k$ singular values with the highest $\Phi(\sigma_{i})$ are retained, and the compressed weight matrix is reconstructed as:
$W_{\text{RLAL}}=\sum_{i=1}^{k}\mathbf{u}_{i}\sigma_{i}\mathbf{v}_{i}^{T}$ (7)
This ensures that $W_{\text{RLAL}}$ preserves the transformations most critical to representation stability, achieving efficient compression while maintaining representation consistency.
@@ 3.4 Layer-wise Low-Rank Decomposition
GRASP applies a layer-wise backward compression strategy, starting from the last redundant layer and moving sequentially backward. For each redundant layer, weight matrices are compressed using Gradient-Guided Singular Value Decomposition, retaining only the top-$k$ singular values ranked by $\Phi(\sigma_{i})$. This ensures a low-rank approximation that preserves representation-critical transformations.
@@ 4 Experiments
In this section, we conduct comprehensive experiments to evaluate GRASP from four key perspectives. (1) We first compare our method with several popular pruning-based LLM compression methods to demonstrate its effectiveness (Section 4.2). (2) Next, we compare GRASP with SVD-based LLM compression methods under different compression ratios with different LLMs (Section 4.3). (3) Then, we further evaluate GRASP on the LongBench (Bai et al., 2023) datasets to assess its performance on long-form text processing tasks (Section 4.4). (4) Finally, we analyze the inference speed-up achieved by GRASP (Section 4.5).
@@ 4.1 Experimental Setup
Below we detail the models, benchmarks and implementation details, with more experimental setups provided in Appendix A.3 due to page limit.
@@
We evaluated our method on a diverse set of LLMs including LLaMA family models LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B, LLaMA2-13B (Touvron et al., 2023b), LLaMA3.1-8B-Instruct (Dubey et al., 2024) and Mistral family models Mistral-7B (Jiang et al., 2023).
Benchmark
10 datasets are used as evaluation benchmarks including two pre-training datasets: WikiText-2 (Merity et al., 2017), PTB (Marcus et al., 1993), seven common sense reasoning datasets include: WinoGrande (Sakaguchi et al., 2020), Hellaswag (Zellers et al., 2019), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-c, ARC-e (Clark et al., 2018), MathQA (Amini et al., 2019) and one generation benchmark GSM8K (Cobbe et al., 2021). Additionally, we include LongBench (Bai et al., 2023) to assess LLM’s performance on long-form text processing after pruning.